OpenStudy (anonymous):

Probability! A teacher has 5 students, 3 boys and 2 girls. She lines them up randomly. How many ways can she line them up? What is the probability of the event that they are lined up with all of the boys before any of the girls (BBBGG)?

OpenStudy (anonymous):

five students, can line up in \(5!\) different ways

OpenStudy (anonymous):

probability that it is BBBGG first one is a boy \[\frac{3}{5}\] second one is a boy givem fist is \[\frac{2}{4}\] third one is a boy\[\frac{1}{3}\] and then surely the last two are girls multiply

OpenStudy (anonymous):

*given the first was

OpenStudy (anonymous):

so if you multiply it do you get 1/10? @satellite73